Exercise · ZW-LAW-OF-TOTAL-PROBABILITY-005 · ★☆☆ Basic

It rains with probability 0.3. If it rains you are late with probability 0.5; otherwise 0.1. What is the probability that you are late?

Choose one option.

Show answer
  • B. 0.22

Explanation

\(0.3\cdot0.5+0.7\cdot0.1=0.15+0.07=0.22\). The 0.15 only counts the rainy case.

Study Probability and statistics in ZWMath →

Review the concept: Law of total probability

The total probability of an event \(A\) is the weighted average of its conditional probabilities \(P(A\mid B_i)\) over a partition \(B_i\) of the sample space, with the probability of each case as its weight.

\[P(A)=\sum_{i=1}^{n}P(A\mid B_i)\,P(B_i)\]

More practice problems: law of total probability

← ZW-LAW-OF-TOTAL-PROBABILITY-004 ZW-LAW-OF-TOTAL-PROBABILITY-006 →